# On finding optimal collective variables for complex systems by minimizing the deviation between effective and full dynamics

### v1

@@ 3 Effective dynamics for a given CV map
In this section, we study the effective dynamics of $X_{n}$ for a given CV map. We give the definition of the effective dynamics in Section 3.1 and then study its properties in Section 3.2. In particular, we show that its transition density can be characterized via a relative entropy minimization to the transition density of $X_{n}$ and discuss the application of this fact in finding CVs (see Proposition 6 and Remark 3.2).
@@ 3.1 Definitions
Before we define the effective dynamics for $X_{n}$, we need to introduce several useful quantities associated to the CV map. To keep the discussions concise and easy to follow, we will adopt the Dirac delta function and work with informal definitions at some places. We point out that all the quantities below can be defined in a rigorous manner (see Remark 3.1).
Let us assume that a $C^{2}$-smooth map $\xi=(\xi_{1},\xi_{2},\dots,\xi_{k})^{\top}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{k}$ is given, where $1\leq k<d$ and $\xi_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ for $i=1,2,\dots,k$. Under certain conditions on $\xi$ (in particular, we assume $\xi$ is a onto map; see [30, 29] for concrete conditions), the level set
$\Sigma_{z}:=\{x\in\mathbb{R}^{d},\xi(x)=z\}\,,$ (38)
@@
for a given $z\in\mathbb{R}^{k}$, is a $(d-k)$-dimensional submanifold in $\mathbb{R}^{d}$ (Figure 1). Given the map $\xi$, the so-called conditional probability measure $\mu_{z}$ on $\Sigma_{z}$ can be defined as
$\mu_{z}(\mathrm{d}{x})=\frac{1}{Q(z)}\pi(x)\delta(\xi(x)-z)\,\mathrm{d}{x}\,,$ (39)
where $\delta(\cdot)$ denotes the Dirac delta function and $Q(z)$ is a normalizing constant given by
$Q(z)=\int_{\mathbb{R}^{d}}\pi(x)\delta(\xi(x)-z)\,\mathrm{d}{x}\,.$ (40)
We also define the measure $\widetilde{\mu}$ on $\mathbb{R}^{k}$ (i.e. the pushforward measure of $\mu$ by $\xi$) by
$\widetilde{\mu}(\mathrm{d}{z})=Q(z)\mathrm{d}{z}\,,\quad z\in\mathbb{R}^{k}\,.$ (41)
The fact that $\widetilde{\mu}$ is a probability measure can be verified using the well-known identity involving the Dirac delta function:
$\int_{\mathbb{R}^{k}}\bigg{[}\int_{\mathbb{R}^{d}}f(x)\delta(\xi(x)-\hat{z})\,\mathrm{d}{x}\bigg{]}\,\mathrm{d}{\hat{z}}=\int_{\mathbb{R}^{d}}f(x)\,\mathrm{d}{x}\,,\quad f:\mathbb{R}^{d}\rightarrow\mathbb{R}~{}\mathrm{is~{}integrable}\,.$ (42)
@@
To motivate the definition of effective dynamics, let us point out that $\xi(X_{n})$ is in general a non-Markovian process in $\mathbb{R}^{k}$, since the probability that $\xi(X_{n+1})\in\widetilde{D}$ at $(n+1)$th step for a subset $\widetilde{D}\subset\mathbb{R}^{k}$ depends on the full state $X_{n}=x$ at $n$th step rather than $\xi(X_{n})=z$ only, i.e.
$\mathbb{P}(\xi(X_{n+1})\in\widetilde{D}|X_{n}=x)=\int_{\widetilde{D}}\int_{\mathbb{R}^{d}}p(x,y)\delta(\xi(y)-\hat{z})\,\mathrm{d}{y}\,\mathrm{d}{\hat{z}}\,.$ (43)
@@
One way to define a transition density in $\mathbb{R}^{k}$ is to take average on the right hand side of (43) over the states $X_{n}=x$ on $\Sigma_{z}$ with respect to the conditional measure $\mu_{z}$ in (39). Specifically, given the CV map $\xi$, we define the effective dynamics of $X_{n}$ as follows.
@@
The effective dynamics of $X_{n}$ associated to the CV map $\xi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{k}$ is defined as the Markov process $Z_{n}$ in $\mathbb{R}^{k}$ with the transition density $\widetilde{p}(z,\hat{z})$, where
$\widetilde{p}(z,\hat{z})=$ $\int_{\Sigma_{z}}\Big{[}\int_{\mathbb{R}^{d}}p(x,y)\delta(\xi(y)-\hat{z})\,\mathrm{d}{y}\Big{]}\,\mu_{z}(\mathrm{d}{x})$ (44)
$=$ $\frac{1}{Q(z)}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}p(x,y)\delta(\xi(x)-z)\delta(\xi(y)-\hat{z})\pi(x)\,\mathrm{d}{x}\,\mathrm{d}{y},\quad z,\hat{z}\in\mathbb{R}^{k}\,.$
Using (42) and the fact that $p(x,\cdot)$ is a probability density in $\mathbb{R}^{d}$, we can verify that $\widetilde{p}(z,\cdot)$ in (44) indeed defines a probability density for $z\in\mathbb{R}^{k}$, i.e. $\int_{\mathbb{R}^{k}}\widetilde{p}(z,\hat{z})\,\mathrm{d}{\hat{z}}=1$. As we show in the next subsection, Definition 44 is very natural, because the invariant distribution of $Z_{n}$ is directly related to the free energy associated to $\xi$ [28] and its transition density $\widetilde{p}$ gives the best approximation of $p$ in the sense of relative entropy minimization [7, Section 2.3].
We conclude this section with a discussion on how to make the definitions above rigorous.
Remark 3.1 (Rigorous definitions).
The conditional probability measure $\mu_{z}$ in (39) and the normalizing constant $Q(z)$ in (40) can be defined rigorously without using the Dirac delta function. See (85) and (86) in Appendix B, respectively. In particular, using co-area formula [12, Theorem 3.11], the transition density in (44) can be written rigorously as
$\widetilde{p}(z,\hat{z})=\int_{\Sigma_{z}}\int_{\Sigma_{\hat{z}}}p(x,y)\big{[}\mathrm{det}(\nabla\xi(y)\nabla\xi(y)^{\top})\big{]}^{-\frac{1}{2}}\,\mu_{z}(\mathrm{d}{x})\,\sigma_{\Sigma_{\hat{z}}}(\mathrm{d}{y}),\quad z,\hat{z}\in\mathbb{R}^{k}\,,$ (45)
where $\sigma_{\Sigma_{\hat{z}}}$ denotes the surface measure on $\Sigma_{\hat{z}}$, $\nabla\xi(y)\in\mathbb{R}^{k\times d}$ denotes the Jacobian matrix of $\xi$ at $y$, and $\mathrm{det}(\nabla\xi(y)\nabla\xi(y)^{\top})$ denotes the determinant of the $k\times k$ matrix $\nabla\xi(y)\nabla\xi(y)^{\top}$.
Figure 1: Illustration of level sets of the CV map $\xi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{k}$ for values $z,\hat{z}\in\mathbb{R}^{k}$.
@@ 3.2 Properties of effective dynamics
To show the invariance of $\widetilde{\mu}$, let us compute, for $\hat{z}\in\mathbb{R}^{k}$,
$\int_{\mathbb{R}^{k}}\widetilde{p}(z,\hat{z})\,Q(z)\,\mathrm{d}{z}$
$=$ $\int_{\mathbb{R}^{k}}\bigg{[}\frac{1}{Q(z)}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}p(x,y)\delta(\xi(x)-z)\delta(\xi(y)-\hat{z})\pi(x)\,\mathrm{d}{x}\,\mathrm{d}{y}\bigg{]}\,Q(z)\,\mathrm{d}{z}$
$=$ $\int_{\mathbb{R}^{k}}\bigg{[}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}p(x,y)\delta(\xi(x)-z)\delta(\xi(y)-\hat{z})\pi(x)\,\mathrm{d}{x}\,\mathrm{d}{y}\bigg{]}\,\mathrm{d}{z}$
$=$ $\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}p(x,y)\delta(\xi(y)-\hat{z})\pi(x)\,\mathrm{d}{x}\,\mathrm{d}{y}$
$=$ $\int_{\mathbb{R}^{d}}\pi(y)\delta(\xi(y)-\hat{z})\,\mathrm{d}{y}$
$=$ $Q(\hat{z})\,,$
where the first equality follows from (44) and (41), the third equality follows from (42), the fourth equality follows from the invariance of $\pi$ in (1), and the last equality follows from the definition of $Q$ in (40). ∎
From now on, we will assume that the effective dynamics $Z_{n}$ is ergodic and $\widetilde{\mu}$ is its unique invariant probability measure. Notice that the free energy associated to the CV map $\xi$, denoted by the map $F:\mathbb{R}^{k}\rightarrow\mathbb{R}$, is related to the normalizing constant in (40) by $F(z)=-\beta^{-1}\ln Q(z)$ [28], where $\beta>0$ is a constant related to the system’s temperature. Therefore, Proposition 5 states that $Z_{n}$ as a surrogate model has the desired invariant probability measure whose density is $Q(z)=\exp(-\beta F(z))$, for $z\in\mathbb{R}^{k}$.
@@
$p(x,y)=p(\hat{z}|x)\,p(y|\hat{z},x),$ (46)
where $p(\hat{z}|x)$ denotes the probability density of $\hat{z}=\xi(y)\in\mathbb{R}^{k}$ given the value $x\in\mathbb{R}^{d}$, and $p(y|\hat{z},x)$ is the probability density of $y\in\Sigma_{\hat{z}}$ given the values $\hat{z}$ and $x$ (see Figure 1 for illustration). Finding a Markov process on $\mathbb{R}^{k}$ that mimics the dynamics of $\xi(X_{n})$ can be achieved by considering the approximation of $p$ within the set $\Theta^{\xi}$ of probability densities $\bar{p}$, which have the following form
$\bar{p}(x,y)=g(\xi(x),\hat{z})\,\bar{p}_{\hat{z}}(y),\quad x,y\in\mathbb{R}^{d},\quad\xi(y)=\hat{z}\in\mathbb{R}^{k}\,,$ (47)
where $g$ is a transition density function on $\mathbb{R}^{k}$ and $\bar{p}_{\hat{z}}$ is a probability density on the level set $\Sigma_{\hat{z}}$ (with respect to its surface measure $\sigma_{\Sigma_{\hat{z}}}$). In other words, we approximate the two terms on the right hand side of (46) by assuming that the first term is only conditioned on $z=\xi(x)$ rather than $x$ and the second term only on $\hat{z}$.
The following result states that, among the probability densities in (47), the one which minimizes the relative entropy (or, KL-divergence) to $p$ is given by the transition density $\widetilde{p}$ (44) of the effective dynamics $Z_{n}$ in Definition 44 and the conditional expectation $\mu_{z}$ (39) (see (85) in Appendix B for the density of $\mu_{z}$ and also Remark 3.1 for discussions).
@@
For $z\in\mathbb{R}^{k}$, let $p_{z}$ be the density of $\mu_{z}$ on $\Sigma_{z}$ with respect to $\sigma_{\Sigma_{z}}$, i.e. $p_{z}=\frac{\mathrm{d}{}\mu_{z}}{\mathrm{d}{}\sigma_{\Sigma_{z}}}$. For fixed $x\in\mathbb{R}^{d}$, denote by $D_{\mathrm{KL}}(p(x,\cdot)\|\bar{p}(x,\cdot))$ the KL-divergence from the density $\bar{p}(x,\cdot)$ to the density $p(x,\cdot)$. Then the minimizer
$\bar{p}^{*}=\arg\min_{\bar{p}\in\Theta^{\xi}}\mathbb{E}_{x\sim\mu}\Big{(}D_{\mathrm{KL}}\big{(}p(x,\cdot)\,\|\,\bar{p}(x,\cdot)\big{)}\Big{)}$ (48)
is given by (47) with $g=\widetilde{p}$ and $\bar{p}_{z}=p_{z}$ for $z\in\mathbb{R}^{k}$.
@@
By definition of KL-divergence [7, Section 2.3], we have, for $\bar{p}\in\Theta^{\xi}$,
$\begin{split}&\mathbb{E}_{x\sim\mu}\Big{(}D_{\mathrm{KL}}\big{(}p(x,\cdot)\,\|\,\bar{p}(x,\cdot)\big{)}\Big{)}\\ =&\int_{\mathbb{R}^{d}}\bigg{[}\int_{\mathbb{R}^{d}}p(x,y)\ln\frac{p(x,y)}{\bar{p}(x,y)}\,\mathrm{d}{y}\bigg{]}\pi(x)\,\mathrm{d}{x}\\ =&\int_{\mathbb{R}^{d}}\bigg{[}\int_{\mathbb{R}^{d}}p(x,y)\ln p(x,y)\,\mathrm{d}{y}\bigg{]}\pi(x)\,\mathrm{d}{x}-\int_{\mathbb{R}^{d}}\bigg{[}\int_{\mathbb{R}^{d}}p(x,y)\ln\bar{p}(x,y)\,\mathrm{d}{y}\bigg{]}\pi(x)\,\mathrm{d}{x}\,,\end{split}$ (49)
@@
where the first term in the last line is independent of $\bar{p}$. For the second term, we can derive
$-\int_{\mathbb{R}^{d}}\bigg{[}\int_{\mathbb{R}^{d}}p(x,y)\ln\bar{p}(x,y)\,\mathrm{d}{y}\bigg{]}\pi(x)\,\mathrm{d}{x}$
$=$ $-\int_{\mathbb{R}^{d}}\bigg{[}\int_{\mathbb{R}^{d}}p(x,y)\big{(}\ln g(\xi(x),\xi(y))+\ln\bar{p}_{\xi(y)}(y)\big{)}\,\mathrm{d}{y}\bigg{]}\pi(x)\,\mathrm{d}{x}$
$=$ $-\int_{\mathbb{R}^{k}}\bigg{[}\int_{\mathbb{R}^{k}}\widetilde{p}(z,\hat{z})\ln g(z,\hat{z})\,\mathrm{d}{\hat{z}}\bigg{]}Q(z)\,\mathrm{d}{z}-\int_{\mathbb{R}^{d}}\bigg{[}\int_{\mathbb{R}^{d}}p(x,y)\ln\bar{p}_{\xi(y)}(y)\,\mathrm{d}{y}\bigg{]}\pi(x)\,\mathrm{d}{x}$
$=$ $-\int_{\mathbb{R}^{k}}\bigg{[}\int_{\mathbb{R}^{k}}\widetilde{p}(z,\hat{z})\ln g(z,\hat{z})\,\mathrm{d}{\hat{z}}\bigg{]}Q(z)\,\mathrm{d}{z}-\int_{\mathbb{R}^{d}}\ln\bar{p}_{\xi(x)}(x)\,\pi(x)\,\mathrm{d}{x}$
$=$ $-\int_{\mathbb{R}^{k}}\bigg{[}\int_{\mathbb{R}^{k}}\widetilde{p}(z,\hat{z})\ln g(z,\hat{z})\,\mathrm{d}{\hat{z}}+\int_{\Sigma_{z}}\ln\bar{p}_{z}(x)\mu_{z}(\mathrm{d}{x})\bigg{]}Q(z)\,\mathrm{d}{z}$
$=$ $\int_{\mathbb{R}^{k}}\bigg{[}D_{\mathrm{KL}}\big{(}\widetilde{p}(z,\cdot)\,\|\,g(z,\cdot)\big{)}+D_{\mathrm{KL}}(p_{z}\,\|\,\bar{p}_{z})\bigg{]}Q(z)\,\mathrm{d}{z}$
$-\int_{\mathbb{R}^{k}}\bigg{[}\int_{\mathbb{R}^{k}}(\widetilde{p}\ln\widetilde{p})(z,\hat{z})\,\mathrm{d}{\hat{z}}+\int_{\Sigma_{z}}\ln p_{z}(x)\mu_{z}(\mathrm{d}{x})\bigg{]}Q(z)\,\mathrm{d}{z}\,,$ (50)

### v2

@@ 3 Effective dynamics for a given CV map
In this section, we study the effective dynamics of $X_{n}$ for a given CV map. We give the definition of the effective dynamics in Section 3.1 and then study its properties in Section 3.2. In particular, we show that its transition density can be characterized via a relative entropy minimization to the transition density of $X_{n}$ and discuss the application of this fact in finding CVs (see Proposition 6 and Remark 3.1). Unless explicitly stated, in this section we do not require the process $X_{n}$ to be reversible.
@@ 3.1 Definitions
Before we define the effective dynamics for $X_{n}$, we need to introduce several quantities associated to the CV map. Given $1\leq k<d$, let us define the space
$\Xi=\Big{\{}\xi\in C^{3}(\mathbb{R}^{d},\mathbb{R}^{k})\,\Big{|}\,\mbox{$\xi$ is onto, $\mbox{rank}(\nabla\xi(x))=k$ at each $x\in\mathbb{R}^{d}$}\Big{\}}\,,$ (36)
where $\nabla\xi(x)\in\mathbb{R}^{d\times k}$ denotes the Jacobian matrix of $\xi$. For $\xi\in\Xi$ and given $z\in\mathbb{R}^{k}$, the level set
$\Sigma_{z}:=\{x\in\mathbb{R}^{d}\,|\,\xi(x)=z\}$ (37)
@@
is a $(d-k)$-dimensional submanifold in $\mathbb{R}^{d}$ (see [39, Remark 1] and [38, Section 1]; see Figure 1 for illustration). We denote by $\sigma_{\Sigma_{z}}$ the surface measure of $\Sigma_{z}$ and we define the probability measure $\mu_{z}$ on $\Sigma_{z}$ as
$\mu_{z}(\mathrm{d}{x})=p_{z}(x)\left[\mathrm{det}(\nabla\xi(x)^{\top}\nabla\xi(x))\right]^{-\frac{1}{2}}\sigma_{\Sigma_{z}}(\mathrm{d}{x})\,,\quad p_{z}(x)=\frac{\pi(x)}{\widetilde{\pi}(z)}\,,$ (38)
where $\widetilde{\pi}(z)$ is the normalizing constant defined by
$\widetilde{\pi}(z)=\int_{\Sigma_{z}}\pi(x)[\mathrm{det}(\nabla\xi(x)^{\top}\nabla\xi(x))]^{-\frac{1}{2}}\,\sigma_{\Sigma_{z}}(\mathrm{d}{x})\,.$ (39)
A determinant term is introduced in (38) and (39), so that by co-area formula (see [18, Theorem 3.11] and [37, Lemma 3.2]) we have, for all integrable functions $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$,
$\int_{\mathbb{R}^{d}}f(x)\mu(\mathrm{d}{x})=\int_{\mathbb{R}^{k}}\Big{(}\int_{\Sigma_{z}}f(x)\,\mu_{z}(\mathrm{d}{x})\Big{)}\widetilde{\mu}(\mathrm{d}{z})\,,$ (40)
where $\widetilde{\mu}$ is the pushforward measure of $\mu$ by the map $\xi$, given by
$\widetilde{\mu}(\mathrm{d}{z})=\widetilde{\pi}(z)\mathrm{d}{z}\,,\quad z\in\mathbb{R}^{k}\,.$ (41)
Note that (40) is identical to the identity in the disintegration theorem for probability measures [1, Theorem 5.3.1]. The measure $\mu_{z}$ is often termed as the conditional probability measure, since by (40) and the law of total expectation it holds
for all integrable functions $f$. In particular, when $\xi$ is a linear map, the level set $\Sigma_{z}$ is a linear subspace and the expressions of quantities introduced above can be largely simplified [34, 35].
@@
To define the effective dynamics for the process $X_{n}$, we borrow the idea from [33] for the effective dynamics of SDEs. Specifically, notice that $\xi(X_{n})$ is in general a non-Markovian process in $\mathbb{R}^{k}$, since the probability that $\xi(X_{n+1})\in\widetilde{D}$ at $(n+1)$th step for a subset $\widetilde{D}\subset\mathbb{R}^{k}$ depends on the full state $X_{n}=x$ at $n$th step rather than $\xi(X_{n})=z$ only, i.e.
$\mathbb{P}(\xi(X_{n+1})\in\widetilde{D}|X_{n}=x)=$ $\int_{\mathbb{R}^{d}}p(x,y)\mathbbm{1}_{\widetilde{D}}(\xi(y))\,\mathrm{d}{y}$ (42)
$=$ $\int_{\widetilde{D}}\left(\int_{\Sigma_{w}}p(x,y)[\mathrm{det}(\nabla\xi(y)^{\top}\nabla\xi(y))]^{-\frac{1}{2}}\sigma_{\Sigma_{z}}(\mathrm{d}{y})\right)\,\mathrm{d}{w}\,,$
@@
where $\mathbbm{1}_{\widetilde{D}}$ denotes the indicator function of the set $\widetilde{D}$ and the expression on the second line follows from co-area formula [37, Lemma 3.2]. Following the idea in [33], one way to define a transition density in $\mathbb{R}^{k}$ is to replace the expression on the right hand side of (42) by its average over the states $X_{n}=x$ on $\Sigma_{z}$ with respect to the conditional measure $\mu_{z}$ in (38). Specifically, given the CV map $\xi$, we define the effective dynamics of $X_{n}$ as follows.
@@
The effective dynamics of $X_{n}$ associated to the CV map $\xi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{k}$ is defined as the Markov process $Z_{n}$ in $\mathbb{R}^{k}$ with the transition density $\widetilde{p}(z,w)$, where
$\widetilde{p}(z,w)=$ $\int_{\Sigma_{z}}\Big{[}\int_{\Sigma_{w}}p(x,y)[\mathrm{det}(\nabla\xi(y)^{\top}\nabla\xi(y))]^{-\frac{1}{2}}\sigma_{\Sigma_{w}}(\mathrm{d}{y})\Big{]}\,\mu_{z}(\mathrm{d}{x})$ (43)
$=$ $\widetilde{\pi}(w)\int_{\Sigma_{z}}\left(\int_{\Sigma_{w}}\frac{p(x,y)}{\pi(y)}\mu_{w}(\mathrm{d}{y})\right)\,\mu_{z}(\mathrm{d}{x})\,,\quad z,w\in\mathbb{R}^{k}\,.$
Note that the second equality in (43) follows from the expression in (38). Using (40) and the fact that $p(x,\cdot)$ is a probability density in $\mathbb{R}^{d}$, we can verify that $\widetilde{p}(z,\cdot)$ in (43) indeed defines a probability density for $z\in\mathbb{R}^{k}$, i.e. $\int_{\mathbb{R}^{k}}\widetilde{p}(z,w)\,\mathrm{d}{w}=1$. As we show in the next subsection, the invariant distribution of $Z_{n}$ is related to the free energy associated to $\xi$ [37, Section 3.2.1] and its transition density $\widetilde{p}$ gives the best approximation of $p$ in the sense of relative entropy minimization [13, Section 2.3]. As discussed above, the effective dynamics defined in Definition 43 can be viewed as the counterpart to the effective dynamics [33] for discrete-in-time and continuous-in-space Markov processes. It also shares similarities with the effective dynamics of Markov chains in previous works [64, 23] and with the asymptotic Markov chains considered in the study of metastability [4, 5, 32].
Figure 1: Illustration of level sets of the CV map $\xi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{k}$ for values $z,w\in\mathbb{R}^{k}$.
@@ 3.2 Properties of effective dynamics
To show the invariance of $\widetilde{\mu}$, let us compute, for $w\in\mathbb{R}^{k}$,
$\left(\int_{\mathbb{R}^{k}}\widetilde{p}(z,w)\widetilde{\mu}(\mathrm{d}{z})\right)\mathrm{d}{w}\,=$ $\widetilde{\pi}(w)\mathrm{d}{w}\int_{\mathbb{R}^{k}}\int_{\Sigma_{z}}\left(\int_{\Sigma_{w}}\frac{p(x,y)}{\pi(y)}\mu_{w}(\mathrm{d}{y})\right)\,\mu_{z}(\mathrm{d}{x})\widetilde{\mu}(\mathrm{d}{z})$
$=$ $\widetilde{\pi}(w)\mathrm{d}{w}\int_{\mathbb{R}^{d}}\left(\int_{\Sigma_{w}}\frac{p(x,y)}{\pi(y)}\mu_{w}(\mathrm{d}{y})\right)\,\mu(\mathrm{d}{x})$
$=$ $\widetilde{\pi}(w)\mathrm{d}{w}\left(\int_{\Sigma_{w}}\frac{\int_{\mathbb{R}^{d}}p(x,y)\mu(\mathrm{d}{x})}{\pi(y)}\,\mu_{w}(\mathrm{d}{y})\right)$
$=$ $\widetilde{\pi}(w)\mathrm{d}{w}=\widetilde{\mu}(\mathrm{d}{w})\,,$
where the first equality follows from the second equality in (43), the second equality follows from (40), the third equality follows from an exchange of integrals, and the fourth equality follows from the invariance of $\mu$ in (1). ∎
From now on, we will assume that the effective dynamics $Z_{n}$ is ergodic and $\widetilde{\mu}$ is its unique invariant probability measure. Notice that, when $X_{n}$ comes from Brownian dynamics (16) with invariant density $\pi(x)=\frac{1}{Z}\mathrm{e}^{-\beta V(x)}$ (Example 2.1), the free energy associated to the CV map $\xi$, denoted by the map $F:\mathbb{R}^{k}\rightarrow\mathbb{R}$, is related to the normalizing constant in (39) by $F(z)=-\beta^{-1}\ln\widetilde{\pi}(z)$ [37, Section 3.2.1], where $\beta>0$ is a constant related to the system’s temperature. Therefore, Proposition 5 states that $Z_{n}$ as a surrogate model has the desired invariant probability measure whose density is $\widetilde{\pi}(z)=\exp(-\beta F(z))$, for $z\in\mathbb{R}^{k}$.
@@
$p(x,y)=p(w|x)\,p(y|w,x),$ (44)
where $p(w|x)$ denotes the probability density of $w=\xi(y)\in\mathbb{R}^{k}$ given the value $x\in\mathbb{R}^{d}$, and $p(y|w,x)$ is the probability density of $y\in\Sigma_{w}$ given the values $w$ and $x$ (see Figure 1 for illustration). Let us define
[…]
which are spaces of probability densities in $\mathbb{R}^{k}$, transition densities in $\mathbb{R}^{k}$, and probability densities on level sets, respectively. One way to find a Markov process on $\mathbb{R}^{k}$ that mimics the dynamics of $\xi(X_{n})$ is to consider the approximation of $p$ within the set $\Theta_{\xi}$ of transition probability densities defined as:
$\begin{split}\Theta_{\xi}=\bigg{\{}g\in C(\mathbb{R}^{d}\times\mathbb{R}^{d},\mathbb{R}^{+})\,\bigg{|}\,&g(x,y)=\widetilde{g}(\xi(x),\xi(y))\,g_{\xi(y)}(y),\,\mbox{for}\,x,y\in\mathbb{R}^{d}\,,\\ &\mbox{where}~{}\widetilde{g}\in\widetilde{\mathcal{G}},~{}(g_{z})_{z\in\mathbb{R}^{k}}\in\mathcal{G}_{\xi,\Sigma}\bigg{\}}\,.\end{split}$ (46)
In other words, we approximate the two terms on the right hand side of (44) by assuming that the first term is only conditioned on $z=\xi(x)$ rather than $x$ and the second term is only conditioned on $w$. Note that the definition of $\mathcal{G}_{\xi,\Sigma}$ in (45) makes sure that, by co-area formula, elements in (46) are transition densities.
The following result states that, among the probability densities in (46), the one which minimizes the relative entropy (or, KL-divergence) to $p$ is given by the transition density $\widetilde{p}$ (43) of the effective dynamics $Z_{n}$ in Definition 43 and the conditional expectation $\mu_{z}$ (38).
@@
For $x\in\mathbb{R}^{d}$, denote by $D_{\mathrm{KL}}(p(x,\cdot)\|g(x,\cdot))$ the KL-divergence from the density $g(x,\cdot)$ to the density $p(x,\cdot)$. Then the minimizer
$g^{\mathrm{opt}}=\arg\min_{g\in\Theta_{\xi}}\mathbb{E}_{x\sim\mu}\Big{(}D_{\mathrm{KL}}\big{(}p(x,\cdot)\,\|\,g(x,\cdot)\big{)}\Big{)}$ (47)
is given by $g^{\mathrm{opt}}(x,y)=\widetilde{p}(\xi(x),\xi(y))\,p_{\xi(y)}(y)$, for $x,y\in\mathbb{R}^{d}$, where $\widetilde{p}$ is the transition density defined in (43) and $p_{z}$ is defined in (38) (in the definition of $\mu_{z}$).
@@
For any $g\in\Theta_{\xi}$ with the expression
$g(x,y)=\widetilde{g}(\xi(x),\xi(y))\,g_{\xi(y)}(y)\,,$ (48)
we show the following identities hold.
$\mathbb{E}_{x\sim\mu}\Big{(}D_{\mathrm{KL}}\big{(}p(x,\cdot)\,\|\,g(x,\cdot)\big{)}\Big{)}-\int_{\mathbb{R}^{d}}\bigg{(}\int_{\mathbb{R}^{d}}p(x,y)\ln p(x,y)\,\mathrm{d}{y}\bigg{)}\mu(\mathrm{d}{x})$
$=$ $-\int_{\mathbb{R}^{d}}\bigg{(}\int_{\mathbb{R}^{d}}p(x,y)\ln\widetilde{g}(\xi(x),\xi(y))\mathrm{d}{y}\bigg{)}\mu(\mathrm{d}{x})-\int_{\mathbb{R}^{d}}\ln g_{\xi(x)}(x)\mu(\mathrm{d}{x})$
$=$ $-\int_{\mathbb{R}^{d}}\bigg{(}\int_{\mathbb{R}^{d}}p(x,y)\ln\widetilde{g}(\xi(x),\xi(y))\mathrm{d}{y}\bigg{)}\mu(\mathrm{d}{x})+\int_{\mathbb{R}^{k}}D_{\mathrm{KL}}(p_{z}\,\|\,g_{z})\widetilde{\mu}(\mathrm{d}{z})$
$-\int_{\mathbb{R}^{d}}\ln p_{\xi(x)}(x)\mu(\mathrm{d}{x})$ (49)
$=$ $\int_{\mathbb{R}^{k}}\bigg{(}D_{\mathrm{KL}}\big{(}\widetilde{p}(z,\cdot)\,\|\,\widetilde{g}(z,\cdot)\big{)}+D_{\mathrm{KL}}(p_{z}\,\|\,g_{z})\bigg{)}\widetilde{\mu}(\mathrm{d}{z})$
$-\int_{\mathbb{R}^{k}}\bigg{(}\int_{\mathbb{R}^{k}}(\widetilde{p}\ln\widetilde{p})(z,w)\,\mathrm{d}{w}\bigg{)}\widetilde{\mu}(\mathrm{d}{z})-\int_{\mathbb{R}^{d}}\ln p_{\xi(x)}(x)\mu(\mathrm{d}{x})\,.$
In the above, $D_{\mathrm{KL}}(p_{z}\,\|\,g_{z})$ denotes the KL divergence of $g_{z}$ from $p_{z}$, both of which are densities on $\Sigma_{z}$ with respect to $\left[\mathrm{det}(\nabla\xi(x)^{\top}\nabla\xi(x))\right]^{-\frac{1}{2}}\sigma_{\Sigma_{z}}(\mathrm{d}{x})$.
@@
We start by calculating the first term on the first line of (49). By definition of KL-divergence [13, Section 2.3], we have, for $g\in\Theta_{\xi}$ with expression in (48),
$\mathbb{E}_{x\sim\mu}\Big{(}D_{\mathrm{KL}}\big{(}p(x,\cdot)\,\|\,g(x,\cdot)\big{)}\Big{)}$ (50)
$=$ $\int_{\mathbb{R}^{d}}\bigg{(}\int_{\mathbb{R}^{d}}p(x,y)\ln\frac{p(x,y)}{g(x,y)}\,\mathrm{d}{y}\bigg{)}\mu(\mathrm{d}{x})$
$=$ $\int_{\mathbb{R}^{d}}\bigg{(}\int_{\mathbb{R}^{d}}p(x,y)\ln p(x,y)\,\mathrm{d}{y}\bigg{)}\mu(\mathrm{d}{x})-\int_{\mathbb{R}^{d}}\bigg{(}\int_{\mathbb{R}^{d}}p(x,y)\ln g(x,y)\,\mathrm{d}{y}\bigg{)}\mu(\mathrm{d}{x})\,,$
@@
where the first term on the last line is independent of $g$. For the second term above, we can derive
$-\int_{\mathbb{R}^{d}}\bigg{(}\int_{\mathbb{R}^{d}}p(x,y)\ln g(x,y)\,\mathrm{d}{y}\bigg{)}\mu(\mathrm{d}{x})$ (51)
=\,= $-\int_{\mathbb{R}^{d}}\bigg{(}\int_{\mathbb{R}^{d}}p(x,y)\left(\ln\widetilde{g}(\xi(x),\xi(y))+\ln g_{\xi(y)}(y)\right)\,\mathrm{d}{y}\bigg{)}\mu(\mathrm{d}{x})$
=\,= $-\int_{\mathbb{R}^{d}}\bigg{(}\int_{\mathbb{R}^{d}}p(x,y)\ln\widetilde{g}(\xi(x),\xi(y))\mathrm{d}{y}\bigg{)}\mu(\mathrm{d}{x})-\int_{\mathbb{R}^{d}}\ln g_{\xi(x)}(x)\mu(\mathrm{d}{x})$
=:absent:=:= : $I_{1}+I_{2}\,,$
where the first equality follows from the expression of $g$ in (48), and the second equality follows from the invariance of $\mu$ in (1). We continue to compute $I_{1}$ and $I_{2}$ above. For $I_{1}$, we have
$I_{1}=$ $-\int_{\mathbb{R}^{d}}\bigg{(}\int_{\mathbb{R}^{d}}p(x,y)\ln\widetilde{g}(\xi(x),\xi(y))\mathrm{d}{y}\bigg{)}\mu(\mathrm{d}{x})$ (52)
$=$ $-\int_{\mathbb{R}^{k}}\bigg{(}\int_{\mathbb{R}^{k}}\widetilde{p}(z,w)\ln\widetilde{g}(z,w)\,\mathrm{d}{w}\bigg{)}\widetilde{\mu}(\mathrm{d}{z})$
$=$ $\int_{\mathbb{R}^{k}}D_{\mathrm{KL}}\big{(}\widetilde{p}(z,\cdot)\,\|\,\widetilde{g}(z,\cdot)\big{)}\widetilde{\mu}(\mathrm{d}{z})-\int_{\mathbb{R}^{k}}\bigg{(}\int_{\mathbb{R}^{k}}(\widetilde{p}\ln\widetilde{p})(z,w)\,\mathrm{d}{w}\bigg{)}\widetilde{\mu}(\mathrm{d}{z})\,.$